OpenStudy (anonymous):

Simplify: 6m + 7n + 5m – 3n A. 11m + 10n B. 15mn C. m + 10n D. 11m + 4n

Vocaloid (vocaloid):

put the m's together 6m + 5m = ?

OpenStudy (anonymous):

11

Vocaloid (vocaloid):

don't forget the m...

OpenStudy (anonymous):

OKay 11m

Vocaloid (vocaloid):

ok, now let's put the n's together... 7n - 3n = ?

OpenStudy (anonymous):

10n

Vocaloid (vocaloid):

7 - 3 = ?

OpenStudy (anonymous):

4

Vocaloid (vocaloid):

yes, so 7n - 3n = ?

OpenStudy (saurabh135):

4n

OpenStudy (saurabh135):

nd hence 11m + 4n

OpenStudy (saurabh135):

nd hence 11m + 4n